Joe borrowed $2500 at 4.5% compounded monthly. After one year he paid $1000 toward the outstanding balance. How much will Joe owe after three years?

(Using Financial calculator method (BAII) by showing inputs within calculator)
Answer #1

Answer

A=p(1+r÷n)nt

Amount after year1=

P=2500

r=4.5

=2500(1+4.5÷n)nt

=2614.85

outstanding after repayment =2614085-1000

=1614.85

outstanding after year3(remaining 2 years) =

p=1614.85

r=4.5

=1614.85(1+4.5÷n)nt

=1766.63